OpenStudy (anonymous):

can someone please help will give medal 25 divided by 5975 show the work please

OpenStudy (anonymous):

25 and under the roof thing is 5975

OpenStudy (prizzyjade):

\[\frac{ 25 }{ 5975 }\]

OpenStudy (anonymous):

i know the answer i did it on my phone but i dont know how to tell them how i got it

OpenStudy (prizzyjade):

\[\frac{ 25 }{ 25 \times 239}\] \[\frac{ 1 }{ 239 }\]